3 men go to buy a watch. Split the cost into 3. The person there says it costs 60. So they pay 20 ea. then the owner comes and says its only 50. Go return 10, he decides to keep four to himself. And gives them ea. 2, but the lost is now 18x3 which is 54. Plus the 4 in his pocket I becomes 58. What happened to the 2? Didn't it equal 60?
1
Expert's answer
2013-01-31T10:58:52-0500
Look 50/3 = 16.67.
Now they are returned 2 rupees each so each got 16.67+2 =18.67.
And 18.67*3 = 56. The other 4 are with the owner. The totalbecomes 60 :)
The reality is that in 50 units, everyone is notsharing 18 or 20, they are sharing 18.67.